OpenStudy (flvs.net):

Medal and Fan Sara left a bin outside in her garden to collect rainwater. She notices that 1/8 gallon of water fills 2/3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words.

OpenStudy (anonymous):

So, One-eighth gallon of water will fill two-thirds of the bin. We want the sentence: "x gallons of water will fill three-thirds of the bin." What we do is multiply two-thirds by its reciprocal (three-halves) and make sure we do it on both sides. \frac{1}{8} = \frac{2}{3} x \frac{3}{2} * \frac{1}{8} = x \frac{3}{16}\ gal. = x So your equation would look like this: \frac{1}{8} * \frac{3}{2}\ gal._{water} = x

OpenStudy (flvs.net):

\[\frac{1}{8} * \frac{3}{2}\ gal._{water} = x\]

OpenStudy (flvs.net):

This?

OpenStudy (anonymous):

yup

OpenStudy (flvs.net):

What is X?

OpenStudy (flvs.net):

What does it equal?

OpenStudy (anonymous):

just use the equation and solve yourself

OpenStudy (flvs.net):

Well what's the answer?

OpenStudy (danjs):

\[\frac{ 1~bin }{ 1 } * \frac{ 1/ 8 ~gal }{ 2/3~bin } = ~~~gal\]

OpenStudy (flvs.net):

I meant

OpenStudy (flvs.net):

How do I solve that @DanJS

OpenStudy (danjs):

you wanted gallons in 1 bin, so start with 1 bin, multiply that by the given ratio of 1/8 gallon to 2/3 bin... notice it is set up so that the units for bin will cancel... just take care of the numbers

OpenStudy (flvs.net):

so I do 1/8 * 2/3?

OpenStudy (danjs):

when dividing a fraction by a fraction, you can flip the bottom one over and multiply it to the top

OpenStudy (flvs.net):

Isn't it "keep Change Flip?"

OpenStudy (danjs):

\[\frac{ 1 }{ 8 }*\frac{ 3 }{ 2 }\]

OpenStudy (flvs.net):

\[\frac{ 1 }{ 8 } * \frac{ 3 }{ 2 } = \frac{ 3 }{ 16 }\] Is that correct?

OpenStudy (danjs):

yep, 3/16 gal per full bin

OpenStudy (flvs.net):

Thank you!